OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = -34 and a9 = -122? -408 -397 -386 -375

OpenStudy (anonymous):

What is the sum of a 52-term arithmetic sequence where the first term is 6 and the last term is 363? 8,487 8,856 9,225 9,594 What is the sum of a 22-term arithmetic sequence where the first term is 57 and the last term is -27? 240 270 300 330 What is the 9th term of the geometric sequence where a1 = -6 and a6 = -6,144? -327,680 -393,216 -458,752 -524,288 What is the 7th term of the geometric sequence where a1 = 128 and a3 = 8? 0.03125 0.06 0.125 0.15625 What is the 6th term of the geometric sequence where a1 = -625 and a2 = 125? -0.2 0.2 -0.04 0.04

OpenStudy (anonymous):

The sum of a 52-term arithmetic sequence where the first term is 6 and the last term is 363 is 9,594. Sum of AP = {n(a + l)} / 2 Where, n = number of terms in Arithmetic progression a = first term of Arithmetic progression l = last term of Arithmetic progression S52 = {n(a + l)} / 2 = {52(6 + 363)} / 2 = 26*369 = 9,594

OpenStudy (anonymous):

3) The 9th term of the geometric sequence where a1 = -7 and a6 = -7,168 is -458,752. a1 = -7 a6 = -7 * r^5 -7,168 = r^5 * (-7) r^5 = -7,168/-7 = 1024 r = 4 The n-th term of a geometric sequence with initial value a and common ratio r is given by an = ar^(n - 1) a9 = -7*4^8 = -458,752

OpenStudy (anonymous):

4) 0.03125

OpenStudy (anonymous):

thank you:)